Operations Report – May 15th

Crew 194 Operations Report 15.05.18

SOL: 2

Name of person filing report: Ben Easter

Non-nominal systems: Power per prior documentation

Notes on non-nominal systems: Continued issues as already described

Generator (hours run): Overnight approximately 10

Solar— SOC 56%

Diesel – 60%

Propane – psi. Not checked

ATV Fuel (Ethanol free gasoline) – 5 gallons.

Water (trailer) – 500 gallons.

Water (static) – Approximately 75%

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 134890

Toilet tank emptied: Yes
Date last emptied: By director prior to our arrival

ATV’s (Honda, 300, 350.1, 350.2, 350.3): all except Honda

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: Approximately 3 gallons

# Hours the ATVs were Used today: 1 hours

Notes on ATVs:

Deimos rover used: No
Hours: N/A
Beginning charge: 100%
Ending charge: XX %
Currently charging: Yes or No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: Yes
Hours: 35.8
Beginning charge: 100 %
Ending charge: 90%
Currently charging: Yes

Opportunity rover used: Yes
Hours: 34.2
Beginning charge: 100 %
Ending charge: 94 %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: XX %
Currently charging: Yes or No

Hab Car used?: No
Where was Hab Car used?:
Why was Hab Car used?:

General notes and comments:

Summary of internet: working well

Summary of suits and radios: working well

Summary of Hab operations: no concerns

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of health and safety issues: None

Questions, concerns and requests to Mission Support: None.

Crew Photos – May 15th

1- Dean Aubin
2- Dani Carroll
3- Dana Levin
4- Phil Jones
5- Sean Serell
6- Doug Krohn
7- Lorna Workman
8- Rich Bishop
9- Poni Bishop
10- Ben Easter

> On May 15, 2018, at 9:27 PM, Jordan Lane wrote: >
>
>
>

Operations Report – May 15th

Crew 194 Operations Report 15.05.18

SOL: 2

Name of person filing report: Ben Easter

Non-nominal systems: Power per prior documentation

Notes on non-nominal systems: Continued issues as already described

Generator (hours run): Overnight approximately 10

Solar— SOC 56%

Diesel – 60%

Propane – psi. Not checked

ATV Fuel (Ethanol free gasoline) – 5 gallons.

Water (trailer) – 500 gallons.

Water (static) – Approximately 75%

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 134890

Toilet tank emptied: Yes
Date last emptied: By director prior to our arrival

ATV’s (Honda, 300, 350.1, 350.2, 350.3): all except Honda

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: Approximately 3 gallons

# Hours the ATVs were Used today: 1 hours

Notes on ATVs:

Deimos rover used: No
Hours: N/A
Beginning charge: 100%
Ending charge: XX %
Currently charging: Yes or No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: Yes
Hours: 35.8
Beginning charge: 100 %
Ending charge: 90%
Currently charging: Yes

Opportunity rover used: Yes
Hours: 34.2
Beginning charge: 100 %
Ending charge: 94 %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: XX %
Currently charging: Yes or No

Hab Car used?: No
Where was Hab Car used?:
Why was Hab Car used?:

General notes and comments:

Summary of internet: working well

Summary of suits and radios: working well

Summary of Hab operations: no concerns

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of health and safety issues: One crew member experienced a medical incident during the EVA related to heat exhaustion while exerting in the heat/sun. The EVA was aborted, and the crew returned to the Hab. The crew member was treated appropriately for heat illness.

Questions, concerns and requests to Mission Support: None.

EVA Report – May 15th

EVA 2 Report

EVA Crew Members – Dani Carroll, Doug Krohn, Lorna Workman, Sean Serell, Phil Jones, Dean Aubin, Rich Bishop, Poni Bishop

Location
Green Mars
Easting: 520500
Northing: 4256600

EVA Objectives
Simulation of medical contingency scenarios during EVA

Results
EVA aborted prior to completion of simulation.

Issues/Mitigation
EVA aborted due to illness in crewmember. Returned to Hab. Crewmember is now feeling well.

ATVs/Rovers Used
ATVs x 4, EVs x2 (Spirit, Phobos)

Sol Summary – May 15th

Crew 194 Daily Summary Report 15.05.2018

MDRS Daily Summary Report for sol 2

Summary Title
Busy day with EVA and Celestial Navigation

Mission Status:
Good

Sol Activity Summary:
Today was Sol 2 and represented our first day waking up in simulation. The weather is quite warm and sunny. We learned about the effects of radiation illness in the morning and performed an EVA involving search and rescue. After the EVA, we studied celestial navigation and using star plots to find locations on a planet. We also welcomed the arrival of crewmember Rick Cole.

Look Ahead Plan: We will continue our learning with discussion of SANS, winter, and water survival training.

Anomalies in work:
EVA aborted due to illness in crewmember.

Weather:Sunny, Very Warm

Crew Physical Status:The affected crewmember has recovered and is feeling well.

EVA: Per above

Reports to be file:
Operations Report
Journalist Report
EVA Report

Support Requested:
Continued anomalies with power system as already documented.

Journalist Report – May 15th

Crew 194 Report
Sol 2

Later than usual start to the morning as most of us surprisingly and coincidentally slept well past sunup. Yesterday evening the crew successfully responded to a simulated medical emergency. With the crew safe and all dangers tended to, we had more energy than usual as the day came to a close. After reviewing the dangers, symptoms, and treatment of radiation exposure to space travelers, we set outside for our EVA. Coincidentally, solar radiation had the landscape and the EVA much warmer than yesterday. Returning from a very successful EVA, we sat down and learned celestial navigation. Enjoying the throwbacks to math, geometry, and using pencil and paper again, we found it pleasantly challenging. Having to say farewell to our instructor Dana, a space nerd with a personality the size of Olympic Mons, we excitedly welcome a new instructor, Rick Cole.

Photo Report – May 15th

Please find attached photos for the Photo Report. Some photos are around the Hab. We did not obtain UTM coordinates for the ones on EVA as we had a crew member become ill (not simulated) and aborted the EVA.

Photo Report – April 24th

Please find attached photos for the Photo Report. Some photos are around the Hab. We did not obtain UTM coordinates for the ones on EVA as we had a crew member become ill (not simulated) and aborted the EVA.